A cylinder of mass 11.0 kg and radius 0.260m rolls without slipping on a horizontal surface. At a particular instant, its center of mass has a speed of 7.30 m/s. Its rotational kinetic energy about its center of mass is then 147J.

a) What is in kg m2 its moment of inertia about its center of mass?

b) What is in J its linear kinetic energy at that instant?

c) What is in J its total kinetic energy at that instant?

d) What is in kg m/s its linear momentum at that instant?

e) What is in kg m2/s its angular momentum about its center of mass at that instant?
Answer #1

Rotational kinetic energy is given as,

E_{R}=0.5I\omega^{2}

we know the relation between linear and rotational velocity as,

v=r\omega\Rightarrow \omega =7.3/0.26=28.07rad/s

(A) Now, moment of inertia I is,

I=2E_{R}/\omega^{2}\Rightarrow I =0.373kg m^{2}

(B) Linear kinetic energy is given as,

K.E.= 0.5mv^2= 293.095J

(C) Total kinetic energy is sum of roational and linear,

Total K.E= 147+239.095=440.095J

(D) Linear momentum is given as,

M= mv= 11*7.3= 80.3kgm/s

(E) Angular momentum is given as,

L=mrv= 11*0.26*7.3=20.88kgm^2/s
Answer #2

1/2*inertia*w^2=147

==>inertia=147*r^2*2/v^2

==>inertia=0.37294 kg m2

Kinetic energy=1/2*mass*v2

                       =11*7.3*7.3/2

                         =293.095 j

Total kinectic enetgy=293.095+147

                                  =440.095 j

Linear momentum=m*v=11*7.3=80.3kgm/s